OpenStudy (lukecrayonz):

Find the magnitude and direction angle of v=-2i+5j

OpenStudy (mertsj):

magnitude = \[\sqrt{(-2)^2+5^2}=\sqrt{29}\]

OpenStudy (lukecrayonz):

I got 5.4 for magnitude and 112 degrees clockwise of the positive x axis.

OpenStudy (mertsj):

That's what I got also. Good job.

OpenStudy (mertsj):

No. Counterclockwise.

OpenStudy (lukecrayonz):

Sorry :P

OpenStudy (mertsj):